OpenStudy (anonymous):

x^? over 2y^? = 1 over 2x^-3 y^4

OpenStudy (callisto):

Consider the right side, \[\frac{1}{2x^{-3} y^4} = \frac{x^3}{2y^4}\] Now, \[\frac{x^?}{2y^?} =\frac{x^3}{2y^4}\] Compare, left side and right side, can you find the answer?

OpenStudy (anonymous):

\[x{3} y ^{4} ?\]

OpenStudy (callisto):

? for x => 3 ? for y =>4